OpenStudy (anonymous):

a cell of bacterium divides into two cells every 20 min. Suppose that there are initially 500 cells. find the number of cells after t hours? how would i solve this?

OpenStudy (anonymous):

in time=0 there is 500 cells and they divdes into two cells in every 20 min so in time t (after t hours) there will be t*60/20 division(i multiplied by 60 to convert it to minute(it was hour)) (for example t is 2hour then what you will do is: 2*60/20=6 really in 2 hour there will be 6 division so my formula is working well) dont forget the formula for division is {{{{{t*60/20}}}}} now you will multply this by initial value i mean 500 and multply result by 2 because they are divided. so what you get is this:500(our initial value)*t*60/20(our formula that i created)*2(because they will be double in each division) so::: 500*t*60/20*2=t*3000